OpenStudy (anonymous):

Given the linear equation x + y = 5, find the slope of its graph. (Medal for best/correct) A. –5 B.–1 C. 1 D. 5

OpenStudy (fedorable):

Write it as y=-x+5 General form y=m*x+b m is the slope im saying the answer is b

OpenStudy (anonymous):

Thanks :) @Fedorable

OpenStudy (fedorable):

no problem, tag me if you need me

OpenStudy (anonymous):

Could you help with one more? Given the linear equation 4y = 8x – 3, find the y-intercept. \[A. -\frac{ 3 }{ 4 }\] \[B. -2\] \[C. \frac{ 3 }{ 4 }\] \[D. 2\]

OpenStudy (anonymous):

@Fedorable ^^ Question Above^^

OpenStudy (fedorable):

im sorry, im stumped with this one ill tag people that can help you @pooja195 @myah @the_ocean_girl @tkhunny @SamsungFanBoy

OpenStudy (anonymous):

That's fine, thanks :)

OpenStudy (fedorable):

@l_o_w_key

OpenStudy (usukidoll):

a linear equation is defined as y = mx+b where m is the slope and b is the y intercept divide everything by 4

OpenStudy (anonymous):

y=2x−3/4 y intercept is -3/4

OpenStudy (usukidoll):

^ ya

OpenStudy (anonymous):

@l_o_w_key Thanks :) I can't give you a medal but i'm going to fan you...

OpenStudy (usukidoll):

-.-

OpenStudy (anonymous):

Same for you @UsukiDoll :P

OpenStudy (anonymous):

lol its okay i don't even understand what medals do..

OpenStudy (usukidoll):

oh thanks XD

OpenStudy (anonymous):

Gonna close this now since my questions were answered :)